What is the theoretical yield of the product (p-sec-butyltoluene) from a Friedel-Craft Alkylation lab.

Chemicals used:

8.0 mL (in one step) + 3.5 mL (in a different step) of toluene (MM: 92.141 g/mol)

0.5 grams of AlCl3 (MM: 133.34 g/mol)

1 mL of HCl (MM: 36.458 g/mol)

3.5 mL of 1-chlorobutane (MM: 92.566 g/mol)

1-2 grams of CaCl2 (MM: 110.98 g/mol).

Product should be p-sec-butyltoluene (MM: 148.249 g/mol)...but p-butyltoluene and o-butyltoluene could form.
Answer #1

Total volume of toluene used = 8.0 + 3.5 mL = 11.5 mL

Mass of toluene used = Volume * Density = 11.5 mL * 0.867 g / mL = 9.97 g

Moles of toluene = 9.97 g / ( 92.141 g/mol ) = 0.108 mol

Volume of 1-chlorobutane = 3.5 mL

Mass of 1- chloro butane = Volume * Density = 3.5 mL * 0.88 g/ mL = 3.08 g

Moles of 1- chlorobutane = 3.08 g / ( 92.566 g/mol ) = 0.033 mol

The Fridel- Crafts reaction occurs at 1:1 mole ratio between toluene and 1- chlorobutane.

So, here the limiting reagent is 1- chlorobutane. Maximum Yield of p-sec-butyltoluene is 0.033 mol

= 0.033 mol * 148.249 g/mol = 4.89 g

Theoretical yield of the product is 4.89 g